# Object-Attribute-Relation Representation based Video Semantic Communication

### v1

@@ III OAR-assisted JSCC and Video Transmission
In this section, video transmission over Gaussian channels is realized, including transmission of reference frames and OAR sequences. OAR-modulated JSCC is proposed for reference frame transmission with reference to DynamicJSCC [12], and OAR sequences is transmitted via LDPC [33] and QAM.
@@ III-A OAR-modulated JSCC
For the coding and transmission of key frames, OAR representation is integrated with Dynamic JSCC[12] as the backbone, consisting of jointly trained source encoder $\mathcal{E}_{\text{se}}$, channel encoder $\mathcal{E}_{\text{ce}}$ and generator $\mathcal{G}$. The overall framework is shown in Fig. 4 (a). The framework contains two parts: OAR feature extraction and OAR-assisted image JSCC.
The OAR layout ${\bf L}_{1}$ is first generated through the layout generation method implemented in section II. Afterwards, deep layout features are obtained through a convolutional layer: ${\bf F}_{\text{l}}=\text{Conv}_{1}({\bf L}_{1})\in\mathbb{R}^{H/4\times W/4\times C}$.
In the backbone of the JSCC networks, OAR modulation module is introduced into the channel encoder $\mathcal{E}_{\text{ce}}$ and generator $\mathcal{G}$, which is shown in Fig. 4 (d). A set of multiplier ${\bf m}\in(0,1)^{H/4\times W/4\times 1}$ and bias ${\bf b}\in(0,1)^{H/4\times W/4\times 1}$ is learned from the feature to be modulated ${\bf F}_{\text{in}}$ and the layout feature ${\bf F}_{\text{l}}$. The feature distribution is shifted by performing a linear operation:
${\bf F}_{\text{out}}={\bf F}_{\text{in}}\odot{\bf m}+{\bf b}$ (9)
@@
$\hat{\bf I}_{1}=\mathcal{G}\left(\mathcal{E}_{\text{ce}}\left(\mathcal{E}_{\text{se}}({\bf I}_{1})\mid{\bf L}_{1}\right)+{\bf n}\mid{\bf L}_{1}\right)$ (10)
With OAR-based modulation, the JSCC system enhances object retention by prioritizing foreground regions. Meanwhile, the integration of OAR features into the generator offers additional semantic cues for the decoding of foreground objects, thereby improving adaptability to higher noise levels.
@@ III-B OAR-based Video Transmission Pipeline
This section integrates the OAR-based video coding system presented in section II with the OAR-assisted image JSCC to propose an OAR-based video transmission framework. Illustrated in Fig. 2, the framework comprises three components: JSCC of key frames, OAR transmission, and OAR-based video reconstruction. OAR-assisted source and channel coding are applied to key frames for transmission via Gaussian channels. Upon reception, reconstruction from the received symbols yields the reconstructed frame $\hat{\bf I}_{1}$, enabling the restoration of non-key frames. OAR transmission is conducted using 1/3-rate LDPC and 4QAM modulation, which realizes distortion-free transmission. With the received reference frames and OAR sequences, the receiver performs video reconstruction based on the video generation network in II-C.
@@ IV-A1 Dataset Preparation
We conduct experiments on the traffic surveillance video dataset UA-DETRAC[34] for performance validation. Given the limitations of current MOT and attribute recognition algorithms in achieving accurate recognition, we utilize the ground truth for subsequent experimental analysis.
Video frames are segmented into a training set and two validation sets with a ratio of 8:1:1 for the number of frames. Validation set val1 shares the same scenes with the training set but features different objects and motions, facilitating validation of generating diverse objects. Conversely, validation set val2 comprises distinct scenes from the training set, enabling assessment of performance across different backgrounds.
@@ IV-A2 Experiment Settings
Models in the proposed framework are trained in two steps. First, the OAR-based video reconstruction model is trained with original reference frames. The loss function, inspired by vid2vid[19], imposes constraints on reconstructed video quality across spatial and temporal dimensions. Subsequently, the OAR-based JSCC network is trained on AWGN channels. Symbol amplitudes are normalized to ensure an average signal power of 1, followed by the addition of noise sampled randomly at SNRs ranging from 0 to 20 dB to the transmitted symbols.
Experiments are conducted to validate performance from two perspectives: video coding and wireless transmission. In both scenarios, video reconstruction is realized via the same model in II-C. However, different compression and transmission methods are employed for reference frames. In the video coding task, reference frames undergo compression via BPG[35] and losslessly transmitted to evaluate the performance of the video reconstruction network. For wireless video transmission, reference frames are transmitted through OAR-assisted JSCC followed by an AWGN channel.
@@ IV-A3 Performance Metrics
We evaluate performance from perspectives of pixel fidelity, perceptual quality, distribution similarity, and task performance. Specifically, PSNR and SSIM assess pixel fidelity, LPIPS[36] represents perceptual quality, while FID[37] and KID[38] measure distribution similarity. Additionally, mAP serves as a metric for task-level evaluation.
PSNR and SSIM metrics, indicative of pixel-level similarity, are computed using TensorFlow. LPIPS, FID and KID metrics extract deep features through neural networks, measuring feature similarity as the metric. LPIPS and FID (KID) are calculated utilizing open-source codes [36] and [39], respectively.
In addition to the aforementioned metrics, this paper also conducts video quality assessment from a task-oriented perspective using mean average precision (mAP) of object detection tasks. An object detection model is trained on the original video frames, and subsequently applied to perform object detection on the reconstructed videos. The mAP metric is then computed to reflect the video quality. In the implementation of this paper, YOLOv5[40] is utilized after fine-tuning.
Figure 5: The performance achieved by the proposed OAR based video coding method with H.265 and DVC at different bit-rates.
In assessing transmission efficiency, three metrics are employed: bit-rate (for lossless transmission), channel bandwidth ratio (CBR, for wireless transmission) and signal-to-noise ratio (for transmission through noisy channels). The bit-rate is calculated by dividing the total number of bits by the video duration. For reference frames in val1 and val2, the average data amount of OAR are 366 bits and 219 bits, respectively. Utilizing a predictive coding approach, the average bit-rates of the OAR sequences are 3.5 kbps and 2.2 kbps, respectively.
CBR is defined as the ratio of the number of transmitted symbols to the original number of source symbols. For dataset val1, an additional CBR of $7.0\times 10^{-4}$ and $2.7\times 10^{-4}$ is expected for the OAR of reference frame and the overall video, respectively. For dataset val2, corresponding CBRs are $4.2\times 10^{-4}$ and $1.7\times 10^{-4}$.
The signal-to-noise ratio $SNR$ is the ratio of signal power $P_{\text{S}}$ to noise power $P_{\text{N}}$. For an AWGN channel $N\sim G(0,\sigma^{2})$, the SNR is expressed as $SNR(dB)=10\log_{10}P_{\text{S}}/\sigma^{2}$. Without loss of generality, $P_{\text{S}}=1$ is adopted, i.e., both the emitting symbols of JSCC and the conventional method are power-normalized.
@@ IV-A4 Baselines and Comparison Algorithms
For the proposed OAR-based video coding system, the performance is compared with conventional video codec H.265 and deep learning-based codec DVC[41]. The H.265 codec utilizes the x265 implementation of ffmpeg[42], and the bit-rate is adjusted by CRFs. DVC is fine-tuned to achieve various bit-rates. Both pretrained model (labeled by pretrained) and the one fine-tuned on UA-DETRAC (labeled by fine-tuned) are examined to mitigate the influence of dataset distribution.
@@
For video transmission over AWGN channels, the experiments consist of transmission of reference frames and transmission of entire videos. For reference frames, the image undergoes encoding with BPG, followed by LDPC[33] coding, and modulation with either BPSK or QAM. BPG implementation is based on the open-source tool[35], while LDPC and BPSK/QAM modulation are adopted from [43]. Three parameter configurations of 1/3 rate (1536,4608) LDPC, 1/2 rate (3072,6144) LDPC and 2/3 rate (3072,4608) LDPC are examined for channel coding, while modulation schemes of BPSK, 4QAM, 16QAM and 64QAM are realized for symbol mapping. In addition, this paper compares the performance of the proposed OAR-based JSCC with the original model to assess the impact of OAR. Both JSCC models utilize identical training methods and network structures to ensure fairness in comparison.
@@
Furthermore, the performance of OAR-based video transmission over AWGN channels is compared with H.265 video coding followed by LDPC and QAM. Besides, the recent video JSCC method DVST[9] is implemented and compared. The DVST model is trained at SNR=10dB and fine-tuned with different rate control parameters to achieve lower CBRs.
IV-B Rate-Distortion Performance of OAR-Based Video Transmission System
@@
Fig. 5 presents the performance comparison of the proposed OAR-based video coding with H.265 and DVC across different metrics on datasets val1 (solid line) and val2 (dashed line). From Fig. 5 (a) and (b), it’s evident that OAR-based video coding is inferior to H.265 and DVC in terms of pixel fidelity. This is because that the proposed coding primarily focuses on downstream vision tasks, where detailed textures of the object and background are either provided by the reference frames or generated directly without rely on residuals. The discarding of prediction residuals plays a crucial role in achieving significant bit-rate reduction.
@@
Despite the absence of residuals, the proposed method outperforms H.265 in terms of feature distribution and perceptual quality. As indicated from Fig. 5 (c) and (d), the proposed method better preserves the distribution properties in the feature dimension, leveraging prior information from trained parameters. Conversely, H.265 exhibits significant feature bias due to the obvious block effect and blurring distortion at low bit-rates. Similarly, according to the LPIPS metrics demonstrated in Fig. 5 (e), the proposed method achieves superior perceptual quality at lower bit-rates (below 100 kbps).
@@
Furthermore, as depicted in Fig. 5 (f), the proposed method maintains high object detection performance as the bit-rate decreases, reflecting the effective reconstruction of foreground objects by OAR. Conversely, H.265 experiences significant performance degradation when the bit-rate falls below 100 kbps. This is because the OAR contains vital object information, enabling better object recovery even at low bit-rates. Specifically, the position and category information within OAR provide object motion and appearance features in the reconstructed video. From another perspective, the OAR-based video coding method achieves 60.56% and 57.08% bit-rate savings when the average mAP reaches 0.7 (val1) or 0.4 (val2). The performance of DVC is inferior to H.265 in general. Furthermore, since DVC and H.265 share identical channel coding and transmission methods, the wireless transmission performance of DVC is not evaluated subsequently.
@@
Fig. 6 displays video frames reconstructed by H.265 and OAR-based coding at a bit-rate of approximately 65 kbps. As shown in the figure, the proposed method significantly reduces the block effect, enhancing perceptual quality compared to H.265 (highlighted in green boxes). Furthermore, objects at edges are better preserved due to the object position and category information in the OAR (marked by purple boxes).
@@
Comparison of val1 and val2 in Figure 5 reveals that methods generally achieve lower performance on val2. This is primarily due to val2 containing more complex motion patterns, posing challenges for inter-frame prediction and compensation. Moreover, val2 contains more complex objects that are inherently harder to detect efficiently, leading to degraded object detection performance across all methods. Moreover, for OAR-based coding, the distribution disparity between val2 and the training set also results in performance losses. As shown in Fig. 6, the object marked by the yellow box suffers from color deviation because the tail of the car is absent in the reference frame, while the background elements marked by the red box gradually fade. Despite this, in practical surveillance video applications, efficient adaptation through model fine-tuning is feasible, given the long processing period and relatively steady backgrounds. To ensure a fair performance comparison and mitigate the impact of the dataset distribution, subsequent analysis predominantly focuses on val2.

### v2

@@ III OAR-assisted JSCC and Video Transmission
In this section, video transmission over Gaussian channels is realized, including transmission of reference frames and OAR sequences. OAR-modulated JSCC is proposed for reference frame transmission with reference to SwinJSCC [42], and OAR sequences is transmitted via LDPC [43] and QAM.
@@ III-A OAR-modulated JSCC
For the coding and transmission of key frames, OAR representation is integrated with SwinJSCC[42] as the backbone, which consists of a jointly trained source encoder $\mathcal{E}_{\text{se}}$, a channel encoder $\mathcal{E}_{\text{ce}}$ and a generator $\mathcal{G}$. The overall framework is shown in Fig. 4 (a), which contains two components: OAR feature extraction and OAR-assisted image JSCC. Note that the backbone network and SNR modulation are derived from the original SwinJSCC[42]. We introduce the proposed OAR modulation in the following.
The OAR layout ${\bf L}_{1}$ is first generated through the layout generation method implemented in section II-C. Deep layout features are subsequently obtained through a Swin Transformer (ST)[44] block: ${\bf F}_{\text{l}}=\text{ST}_{1}({\bf L}_{1})\in\mathbb{R}^{H/4\times W/4\times C}$.
In the backbone of the JSCC networks, the OAR modulation module is introduced into the channel encoder $\mathcal{E}_{\text{ce}}$ and generator $\mathcal{G}$, which is shown in Fig. 4 (c). A set of multipliers ${\bf m}\in(0,1)^{h_{\text{in}}\times w_{\text{in}}\times 1}$ is learned from the layout feature ${\bf F}_{\text{l}}$ utilizing three consecutive ST blocks. The feature distribution is shifted by performing a pointwise multiplication operation:
${\bf F}_{\text{out}}={\bf F}_{\text{in}}\odot{\bf m}.$ (10)
@@
$\hat{\bf I}_{1}=\mathcal{G}\left(\mathcal{E}_{\text{ce}}\left(\mathcal{E}_{\text{se}}({\bf I}_{1})\mid{\bf L}_{1}\right)+{\bf n}\mid{\bf L}_{1}\right).$ (11)
With OAR-based modulation, the JSCC system enhances object retention by prioritizing foreground regions. Moreover, the integration of OAR features into the generator offers additional semantic cues for the decoding of foreground objects, thereby improving adaptability to higher noise levels.
@@ III-B OAR-based Video Transmission Pipeline
This section integrates the OAR-based video coding system presented in section II with the OAR-modulated image JSCC to propose an OAR-based video transmission framework. As illustrated in Fig. 2, the framework comprises three components: JSCC of key frames, OAR transmission, and OAR-based video reconstruction. OAR-assisted source and channel coding are applied to key frames for transmission via Gaussian channels. Upon reception, reconstruction from the received symbols yields the reconstructed frame $\hat{\bf I}_{1}$, enabling the restoration of subsequent non-key frames. OAR transmission is conducted via 1/3-rate LDPC and 4QAM modulation, which achieves distortion-free transmission. With the received reference frames and OAR sequences, the receiver performs video reconstruction based on the video generation network in II-C.
@@ IV Experiments and Results
Figure 5: Performance comparison of the proposed OAR based method with H.264, H.265 and DVC at different bit-rates on the UA-DETRAC dataset.
@@ IV-A1 Dataset Preparation
We conduct experiments mainly on the traffic surveillance video dataset UA-DETRAC[45] for performance validation, and the videos are resized to $512\times 512$. Given the limitations of current MOT and attribute recognition algorithms in achieving accurate recognition, we utilize the ground truth for subsequent experimental analysis. Additionally, we conducted experiments on the CATER[46] and SoccerNet[47] datasets to demonstrate the generalizability of the proposed framework to other scenarios. The CATER dataset consists of videos of virtual scenes generated via 3D rendering. Objects such as spheres, squares, cones and cylinders of different sizes, colors and materials perform actions such as translations and rotations. The dataset is created for future immersive scene understanding, and experiments on this dataset help to validate the effectiveness of the proposed video transmission method for future immersive media and videos of interactive scenes. The SoccerNet dataset includes videos of several real soccer games, where the locations of players and balls are annotated. The videos of CATER and SoccerNet are resized to $256\times 256$ and $512\times 512$, respectively.
For dataset UA-DETRAC, video frames are segmented into a training set and two validation sets with a ratio of 8:1:1 for the number of frames. The validation set val1 shares the same scenes as the training set but features different objects and motions, facilitating the validation of generating diverse objects. This is adapted to scenarios where the model is deployed on a fixed camera and trained with historical data from that camera. Conversely, the validation set val2 comprises scenes distinct from those of the training set, enabling an assessment of performance across different backgrounds. For the CATER and SoccerNet datasets, the training and validation sets are divided at the ratio of 4:1 and taken from videos of different scenes.
@@ IV-A2 Experiment Settings
The models in the proposed framework are trained in two steps. First, the OAR-based video reconstruction model is trained with original reference frames. The optimization object shown in expression (9) imposes constraints on reconstructed video quality across spatial and temporal dimensions as well as optical flow estimation accuracy. The OAR-based JSCC network is distinct trained on AWGN channels. The symbol amplitudes are normalized to ensure an average signal power of 1, followed by the addition of AWGN noise sampled randomly at SNRs ranging from 0 to 20 dB to the transmitted symbols.
Experiments are conducted to validate the performance from two perspectives: video coding and wireless transmission. In both scenarios, video reconstruction is realized via the same model in II-C. However, different compression and transmission methods are employed for reference frames. In the video coding task, reference frames undergo compression via BPG[48] and are losslessly transmitted to evaluate the performance of the video reconstruction network. For wireless video transmission, reference frames are transmitted through OAR-assisted JSCC followed by an AWGN channel.
@@ IV-A3 Performance Metrics
We evaluate performance from the perspectives of pixel fidelity, perceptual quality, distribution similarity, and task performance. Specifically, PSNR and SSIM assess pixel fidelity, LPIPS[49] represents perceptual quality, and FID[50] and KID[51] measure distribution similarity. Additionally, mAP serves as a metric for task-level evaluation.
The PSNR and SSIM metrics, which are indicative of pixel-level similarity, are computed via TensorFlow. LPIPS, FID and KID metrics extract deep features through neural networks, measuring feature similarity as the metric. LPIPS and FID (KID) are calculated via open-source codes [49] and [52], respectively.
In addition to the aforementioned metrics, this paper also conducts video quality assessment from a task-oriented perspective using the mean average precision (mAP) of object detection tasks. An object detection model is trained on the original video frames, and subsequently applied to perform object detection on the reconstructed videos. The mAP metric (mAP@.95) is then computed to reflect the video quality. In the implementation of this paper, YOLOv5[53] is utilized after fine-tuning.
To assess transmission efficiency, three metrics are employed: the bit-rate (for lossless transmission), the channel bandwidth ratio (CBR, for wireless transmission) and the signal-to-noise ratio (SNR, for transmission through noisy channels). The bit-rate is calculated by dividing the total number of bits by the video duration. For reference frames in val1 and val2 of UA-DETRAC, the average data amounts of OAR are 366 bits and 219 bits, respectively. Using a predictive coding approach, the average bit-rates of the OAR sequences are 3.5 kbps and 2.2 kbps, respectively. For the CATER and SoccerNet datasets, the average bit-rate of the OAR sequences are 3.28 kbps and 4.49 kbps, respectively.
CBR is defined as the ratio of the number of transmitted symbols to the original number of source symbols. For dataset val1 of UA-DETRAC, additional CBRs of $7.0\times 10^{-4}$ and $2.7\times 10^{-4}$ are expected for the OAR of the reference frame and the overall video, respectively. For val2, the corresponding CBRs are $4.2\times 10^{-4}$ and $1.7\times 10^{-4}$. The average CBRs of OAR sequences for the CATER and SoccerNet dataset are $1.02\times 10^{-3}$ and $1.40\times 10^{-3}$, respectively. The variation in the average CBR of OAR sequences on different datasets is caused mainly by the resolution and object density.
The signal-to-noise ratio (SNR) is the ratio of the signal power $P_{\text{S}}$ to the noise power $P_{\text{N}}$. For an AWGN channel $N\sim G(0,\sigma^{2})$, the SNR is expressed as $SNR(dB)=10\log_{10}P_{\text{S}}/\sigma^{2}$. Without loss of generality, $P_{\text{S}}=1$ is adopted, i.e., both the emitting symbols of JSCC and the conventional method are power-normalized.
@@ IV-A4 Baselines and Comparison Algorithms
For the proposed OAR-based video coding system, the performance is compared with conventional video codec H.265 and the deep learning-based codec DVC[54]. The H.265 codec utilizes the x265 implementation of ffmpeg[55], and the bit-rate is adjusted by CRFs. DVC is fine-tuned to achieve various bit-rates. Both the pretrained model (labeled by pretrained) and the one fine-tuned on specific datasets (labeled by fine-tuned) are examined to mitigate the influence of dataset distribution.
@@
For video transmission over AWGN channels, the experiments involve the transmission of reference frames and transmission of entire videos. For reference frames, the image undergoes encoding with BPG, followed by LDPC[43] coding, and modulation with either BPSK or QAM. BPG implementation is based on the open-source tool[48], and LDPC and BPSK/QAM modulations are adopted from [56]. Three parameter configurations of 1/3 rate (1536,4608) LDPC, 1/2 rate (3072,6144) LDPC and 2/3 rate (3072,4608) LDPC are examined for channel coding, and modulation schemes of BPSK, 4QAM, 16QAM and 64QAM are utilized for symbol mapping. In addition, this paper compares the performance of the proposed OAR-based JSCC with the original model to assess the impact of OAR. Both JSCC models utilize identical training methods and network structures to ensure fairness in comparison.
@@
Furthermore, the performance of OAR-based video transmission over AWGN channels is compared with H.265 video coding followed by LDPC and QAM. In addition, the recent video JSCC method DVST[9] is implemented and compared. The DVST model is trained at SNR=10 dB and fine-tuned with different rate control parameters to achieve lower CBRs.
Figure 7: Performance comparison of the proposed OAR based method with H.264, H.265 and DVC at different bit-rates on the CATER dataset.
Figure 8: Performance comparison of the proposed OAR based method with H.265 and DVC at different bit-rates on the SoccerNet dataset.
IV-B Rate-Distortion Performance of OAR-Based Video Coding System
@@
Fig. 5 presents the performance comparison of the proposed OAR-based video coding framework with H.265 and DVC across different metrics on the UA-DETRAC dataset. The solid and dashed lines represent the performance on val1 and val2, respectively. From Fig. 5 (a) and (b), it is evident that OAR-based video coding is inferior to H.265 and DVC in terms of pixel fidelity. This is because the proposed coding focuses primarily on downstream vision tasks, where detailed textures of the object and background are either provided by the reference frames or generated directly without relying on residuals. The discarding of prediction residuals plays a crucial role in achieving significant bit-rate reduction.
@@
Despite the absence of residuals, the proposed method outperforms H.265 in terms of feature distribution and perceptual quality. As indicated in Fig. 5 (c) and (d), the proposed method better preserves the distribution properties in the feature dimension. Conversely, H.265 exhibits significant feature bias due to the obvious block effect and blurring distortion at low bit-rates. Similarly, according to the LPIPS metrics demonstrated in Fig. 5 (e), the proposed method achieves superior perceptual quality at lower bit-rates (especially below 100 kbps).
@@
Furthermore, as depicted in Fig. 5 (f), the proposed method maintains high object detection performance as the bit-rate decreases, reflecting the effective reconstruction of foreground objects by OAR. Conversely, H.265 experiences significant performance degradation when the bit-rate falls below 100 kbps. This is because the OAR contains vital object semantics, enabling better object recovery even at low bit-rates. Specifically, the position and category information within the OAR provides object motion and appearance features in the reconstructed video. From another perspective, the OAR-based video coding method achieves 60.56% and 57.08% bit-rate savings when the average mAP reaches 0.7 (val1) or 0.4 (val2), respectively. The performance of DVC is inferior to H.265 in general.
@@
Fig. 6 displays video frames reconstructed by H.265 and OAR-based coding at a bit-rate of approximately 65 kbps. As shown in the figure, the proposed method significantly reduces the block effect, enhancing the perceptual quality compared with H.265 (highlighted in the green boxes). Furthermore, objects at edges are better preserved because of the object position and category information in the OAR (marked by purple boxes).
Figure 9: The image perceptual quality (LPIPS) at different CBRs and SNRs on the UA-DETRAC dataset. (a)-(b) illustrate the LPIPS versus SNR at different CBRs, and (c)-(d) illustrate the LPIPS versus CBR at different SNRs.
Figure 10: Performance comparison of OAR-based video transmission to H.265 paired with LDPC and DVST on the UA-DETRAC dataset. (a)-(b) are performances at CBR of 1/300, and (c)-(d) at SNR of 0 dB. The H.265 and H.264-based transmission under some parameters suffer from decoding failure and are not plotted. Same with Fig. 11 and Fig. 12
@@
A comparison of val1 and val2 in Fig. 5 reveals that methods generally achieve lower performance on val2. This is primarily due to val2 containing more complex motion patterns, posing challenges for inter-frame prediction and compensation. Moreover, val2 contains more complex objects that are inherently harder to detect efficiently, leading to degraded object detection performance across all methods. For OAR-based coding, the distribution disparity between val2 and the training set also results in performance degradation. As shown in Fig. 6, the object marked by the yellow box suffers from color deviation because the tail of the car is absent in the reference frame, whereas the background elements marked by the red box gradually fade. Despite this, in practical surveillance video applications, efficient adaptation through model fine-tuning is feasible, given the long processing period and relatively steady backgrounds. To ensure a fair performance comparison and mitigate the impact of the dataset distribution, subsequent analysis predominantly focuses on val2.
The video coding performances of the different methods on CATER and SoccerNet datasets are plotted in Fig. 7 and Fig. 8, respectively. For brevity, only four metrics, PSNR, FID, LPIPS and mAP, are presented. H.265 suffers from severe performance degradation due to more overheads on the low-resolution CATER dataset. Therefore, we also present the performance of H.264 to more fully demonstrate the advantages of the proposed method over conventional video coding frameworks. A comparison of Fig. 7 with Fig. 5 (val2) reveals that the proposed method achieves higher performance on the CATER dataset in all the metrics. This is because the CATER dataset has a relatively homogeneous background and the objects do not have complex textures, which is favorable for more fidelity video generative reconstruction. Fig. 7(b) and (c) demonstrate that the proposed method achieves higher perceptual quality compared to H.265 and DVC, which is consistent with the experimental results on the UA-DETRAC dataset. In Fig. 7(d), the fine-tuned DVC achieves comparable performance to the proposed method, which is due to the homogeneous distribution characteristics of the dataset. As observed in Fig. 8 (d), the performance of the proposed method in object detection on the SoccerNet dataset does not suffer from significant degradation with decreasing bit-rate. This results from the fact that there are only two categories in the dataset: person and ball. The generative model has a more explicit objective, and thus enables semantically consistent object generation despite the loss of texture information (at very low bit-rates) in the reference frame.
Figure 11: Performance comparison of the proposed OAR based method with H.264 and H.265 paired with LDPC and DVST on the CATER dataset.
Figure 12: Performance comparison of the proposed OAR based method with H.265 paired with LDPC and DVST on the SoccerNet dataset.